OpenStudy (anonymous):

find the value of 8w-9 given that -3w+8=5..simplify as much as possible: 8w-9=

OpenStudy (aroub):

wait.. u want the value of -3w+8=5 and then the answer substitute it by 8w-9 ??

OpenStudy (anonymous):

should be 8w-9 given that -3w+8=5

OpenStudy (aroub):

okay.... as i understood -3w+8=5 -3w=-8+5 -3w=-3 w=1 8w-9= 8(1)-9= 8-9=-1

OpenStudy (anonymous):

Thanks